a b and c working together complete a job in 18 days. A and B together work twice as much as C; A and C together work thrice as much as B. A alone can finish the work in ?
Answer by richwmiller(17219)

1/a+1/b+1/c=1/18,
a+b=2c,
a+c=3b
a = 141/2, b = 423/10, c = 282/5
a=70.5, b=42.3, c=56.4